OpenStudy (anonymous):

The mass of an electron is 9.11× 10^–31 kg. What is the uncertainty in the position of an electron moving at 6.00 × 10^6 m/s with an uncertainty of Δv = 0.01 × 10^6 m/s?

OpenStudy (jfraser):

plug into the uncertainty equation:\[\Delta x * \Delta m*v \ge \frac{h}{4\pi}\]